# MAD Speech: Measures of Acoustic Diversity of Speech

### v1

@@ Abstract
Generative spoken language models produce speech in a wide range of voices, prosody, and recording conditions, seemingly approaching the diversity of natural speech. However, the extent to which generated speech is acoustically diverse remains unclear due to a lack of appropriate metrics. We address this gap by developing lightweight metrics of acoustic diversity, which we collectively refer to as MAD Speech. We focus on measuring five facets of acoustic diversity: voice, gender, emotion, accent, and background noise. We construct the metrics as a composition of specialized, per-facet embedding models and an aggregation function that measures diversity within the embedding space. Next, we build a series of datasets with a priori known diversity preferences for each facet. Using these datasets, we demonstrate that our proposed metrics achieve a stronger agreement with the ground-truth diversity than baselines. Finally, we showcase the applicability of our proposed metrics across several real-life evaluation scenarios. MAD Speech will be made publicly accessible.
@@ 1 Introduction
These generative models are typically evaluated in terms of faithfulness to their inputs, i.e. speaker identity and transcript (Borsos et al., 2023a; Wang et al., 2023; Kharitonov et al., 2023). Thus it remains unknown how diverse their outputs are, since those metrics neither capture variability nor take into account factors of the perceived speech diversity (i.e., recording conditions, prosody, styles, and accents). Being able to measure speech diversity would also help detect mode collapse when finetuning models (Kirk et al., 2023), selecting training & inference hyperparameters, and optimizing for human feedback (Cideron et al., 2024) Moreover, such measurements are vital when building synthetic data or mixing existing natural datasets.
@@
The suite of MAD Speech metrics will be made publicly accessible.
@@ 2 Background
Different diversity metrics have been proposed for evaluating generative models in computer vision (Heusel et al., 2017; Naeem et al., 2020; Sajjadi et al., 2018; Kynkäänniemi et al., 2019) and natural language processing (Zhang et al., 2020; Li et al., 2016; Zhu et al., 2018; Shen et al., 2019; Dušek et al., 2019; Shu et al., 2019). However these metrics are either computed on raw inputs for text (Zhu et al., 2018; Shen et al., 2019; Alihosseini et al., 2019) or based on underlying representations specific to text and images (Heusel et al., 2017; Cífka et al., 2018; Tevet and Berant, 2021) which makes them unsuitable for evaluating acoustic diversity in generative speech models.
@@ 3 Learning Representations
We take a two-steps approach for measuring diversity of a set of speech utterances. First, we encode raw waveform audio in a compact representation and, afterwards, we estimate the diversity of a set of utterances in this embedding space. Hence, it is vital that the information about the relevant properties of speech is preserved in the embedding space, while the irrelevant details are omitted. We calculate embeddings in a hierarchical way: at first, we map audio to general-purpose embeddings and then apply lightweight specialized projection models that are tailored to capture a specific facet. As a result, our approach allows us to get reliable per-facet diversity metrics.
SpeechSim We aim to learn speech representations where acoustically similar audio segments are close to each other in the embedding space. Following Saeed et al. (2021), we train a self-supervised contrastive model with positive pairs coming from non-overlapping fixed-length chunks of the same speech utterance. The main idea being that such speech segments are likely to be acoustically similar while being semantically different — as there is no overlap between positive pairs. We refer to this model as SpeechSim. SpeechSim provides a speech representation that are sensitive to changes in acoustic diversity, however, it has all types of acoustic properties mixed up together. Indeed, our experiments in Section 7 show that the vanilla SpeechSim embeddings are strongly biased towards speaker voice. We therefore learn projection models (heads) on top of SpeechSim embeddings to extract different types of acoustic information.
@@ 5.1 Our Approach
In the following, we describe the general template for building a series of datasets of controllable diversity, using voice diversity as a reference. We fix the sample size a priori and we define voice diversity as the number of distinct speakers in a set of examples. To build one dataset, we sample examples from a natural dataset such that we obtain the required number of speakers and have every speaker represented by the same number of utterances. We create a series of datasets by varying the number of distinct speakers. We also ensure that other facets do not differ from one dataset to another as much as possible so that we make sure we indeed test speaker voice diversity.
@@ 5.2 Evaluation data
We build evaluation (validation) sets from LibriTTS (Zen et al., 2019) test-clean & test-other (dev-clean & dev-other). Each set contains 200 utterances of a single gender (either all male voices or all female voices) to test whether a particular metric is sensitive to the voice diversity independently from the gender diversity. We vary the number of distinct voices in $\{5,10,15,20,25,33\}$. For each diversity level, we randomly sample 100 sets using different seeds.
@@
We build evaluation (validation) sets from test (dev) sets of VCTK (Veaux et al., 2016). Accent classes are common between train, dev, and tests sets. Again, we built two types of samples (1) with equal number of different speaker voices and (2) with a unique speaker voice for high-entropy datasets but all speaker voices— equally represented—for low-entropy datasets. Speakers in test and dev sets are distinct from those found in the training set.
@@
We build evaluation and validation sets by sampling 20% and 10% of AudioSet, respectively (Gemmeke et al., 2017). We select examples that have exactly two tags, one of them being “Speech”. We treat the second tag as a class of the background noise. Next, we construct a sample by we randomly selecting 100 examples with varying number of distinct noise classes (1, 5, 10, 25, 50, or 100).
@@
6 Experimental Setup
6.1 Data
6.2 Training details
SpeechSim
Projection models
We train projection models on top of SpeechSim embeddings in a contrastive manner. The positive pairs are constructed by taking two examples with the same labels (e.g., same emotion). Except for the gender projector, all models have 2 fully connected layer (of size 256 and 128) with a GELU activation (Hendrycks and Gimpel, 2016) in-between. In the case of the gender projector, we found that a 4-layers network works better (dimensions of 256, 256, 128 and 128). Dropout rate of 0.1 is applied to SpeechSim input representations. We trained the projection models using Adam (Kingma and Ba, 2014) and use a learning rate of 1e-4, and a batch size of 128. We also apply a weight decay of 1e-3 for projection models related to speaker voices, emotion and background noise and 1e-4 for other facets. We select hyperparameters using validation loss.
6.3 Baseline Representations
Wav2Vec-BERT
HuBERT
SoundStream
@@ 6.3 Baseline Representations
Training from scratch
@@ 7.1 Variability in a single facet
From Table 1 we firstly notice that, across all columns, SpeechSim shows higher correlation scores than off-the-shelf embedding models and the models trained from scratch. This supports our decision to rely on SpeechSim as a basis for a per-facet models. In turn, further voice specialization of SpeechSim (SpeechSim/Voice) results in similar scores. Finally, on comparing average cosine vs. Vendi Score aggregation, we observe that often Vendi Score shows higher correlations.
@@
In Table 3 we report average Spearman correlations for the tested representation models, when tested on EmoV and Expresso separately. From these results, we see that, generally, SpeechSim performs better than other general-purpose representations across all configurations. Equally, the specialized projection of SpeechSim for the emotion facet, SpeechSim/Emotion performs best across all representations. As with voice diversity, Vendi Score gets higher correlations than average cosine dissimilarity.
To make sure that our models can measure diversity of emotions beyond that is covered by the (limited) label sets of EmoV and Expresso, we run an experiment where the projection model is trained on one dataset and tested on another (the only common label is “neutral”). Table 3 shows that SpeechSim finetuned on Expresso does transfer to EmoV classes as it obtains a high Spearman correlation. However, SpeechSim finetuned on EmoV gets a Spearman correlation equal to 0.321. Digging deeper, we additionally compute the correlation of the SpeechSim/Emotion-EmoV scores averaged across seeds instead of the average correlation of all scores as computed in Table 3. In this, SpeechSim/Emotion-Emov gets a correlation of 0.964 which means that SpeechSim finetuned on EmoV generalizes well on average.
@@
We report our results in Table 5. We see that generally all representations have somewhat low levels of correlation. SpeechSim/Noise combined with Vendi Score obtains the second highest correlation, following HuBERT.
We additionally compute the correlation of the SpeechSim/Noise scores averaged for all seeds instead of the average correlation of all scores as computed in Table 5. In this setup, SpeechSim/Noise gets a Spearman correlation of 1.00, indicating that it works very well in average.
@@ 7.2 Are facets mixed within metrics?
Emotions and Voices
@@
We run a similar experiment, examining interactions between emotion and voice diversity within metrics. This time, we change the number of speakers to be 1 or 4. We report the results in Table 7. We observe that SpeechSim/Emotion is unaffected by the changes in the number of speakers, while SpeechSim/Voice follows the speaker diversity and hence it has a negative correlation with the emotion diversity score. SpeechSim focuses on the changes in the speaker diversity, thus having a negative correlation with the gender diversity.
Accents and Voices
@@
In this experiment, we use 7 and 30 voices. The results are reported in Table 8. We see that the projected metric is unaffected by changes in voices, while the vanilla version of SpeechSim is focused on the speaker facet.
@@
Overall, the experiments reported in this Section support the necessity of introducing per-facet projection models. While SpeechSim embeddings might provide reasonably good correlations with the ground-truth diversity when only one facet is changed, they implicitly mix all facets into a single score. We find this inconvenient, as (a) changes in one facet can mask changes in another, (b) the user might want to control the trade-off explicitly. Hence, per-facet models are preferrable.
8 Comparing Generative Models
We start by assessing whether the introduction of SoundStorm (Borsos et al., 2023b) to the AudioLM framework Borsos et al. (2023a) lead to any hidden penalties in acoustic diversity (Section 8.1). Next, we examine the effect of the Best-K sampling technique used by Kharitonov et al. (2023) on acoustic diversity (Section 8.2). We study how temperature changes acoustic diversity of the synthesized speech (Section 8.3) and compare acoustic diversity of off-the-shelf TTS models (Section 8.4).
8.1 Semantic-to-Acoustic Token Conversion: SoundStorm vs. AudioLM
In Table 9 we report, for each facet of the acoustic diversity, the ratio of utterances where AudioLM generated higher-diversity audio than SoundStorm. We see that SoundStorm produces audio that is more diverse in voices, emotions, accents, and background noise, while being behind is the gender diversity. We hypothesize SoundStorm managed to exploit speaker details that “leak” in the semantic tokens, hence biasing the produced outputs.
Voice Gender Emotion Accent Background noise
0.0 62.5 6.3 1.6 4.7
Table 9: Semantic-to-acoustic tokens resynthesis task. The proportion of utterances (%) where the AudioLM stage 2 model generates samples that are more diverse than those generated by SoundStorm.
8.2 SPEAR-TTS: Best-of-K decoding
K Voice Gender Emotion Accent Background Noise
1 64 41 62 63 61
4 0 27 0 0 0
8.3 Text Synthesis: Influence of Temperature
T Voice Gender Emotion Accent Background Noise
0.7 12 33 21 18 29
0.9 45 21 40 40 36
Table 11: $T\in\{0.7,0.9\}$ have higher diversity scores than with $T=0.8$.
@@ 8.3 Text Synthesis: Influence of Temperature
8.4 Acoustic diversity in Text-to-Speech systems
Next, we measure acoustic diversity of several publicly available TTS systems. Here, our goal is to show that MAD Speech brings a new dimension to the evaluation of real-world TTS systems. We compare the following systems:

### v2

@@ Abstract
Generative spoken language models produce speech in a wide range of voices, prosody, and recording conditions, seemingly approaching the diversity of natural speech. However, the extent to which generated speech is acoustically diverse remains unclear due to a lack of appropriate metrics. We address this gap by developing lightweight metrics of acoustic diversity, which we collectively refer to as MAD Speech. We focus on measuring five facets of acoustic diversity: voice, gender, emotion, accent, and background noise. We construct the metrics as a composition of specialized, per-facet embedding models and an aggregation function that measures diversity within the embedding space. Next, we build a series of datasets with a priori known diversity preferences for each facet. Using these datasets, we demonstrate that our proposed metrics achieve a stronger agreement with the ground-truth diversity than baselines. Finally, we showcase the applicability of our proposed metrics across several real-life evaluation scenarios. MAD Speech is made publicly available [^1] https://github.com/MatthieuFP/MAD_Speech.
MAD Speech: Measures of Acoustic Diversity of Speech
Matthieu Futeralthanks:   Work done while interning at Google DeepMind. Correspondence to matthieu.futeral@inria.fr Inria Paris                        Andrea Agostinelli Google DeepMind                        Marco Tagliasacchi Google DeepMind
Neil Zeghidour Kyutai                        Eugene Kharitonovthanks:   Correspondence to kharitonov@google.com Google DeepMind
@@ 1 Introduction
These generative models are typically evaluated in terms of faithfulness to their inputs, i.e. speaker identity and transcript (Borsos et al., 2023a; Wang et al., 2023; Kharitonov et al., 2023). Thus it remains unknown how diverse their outputs are, since those metrics neither capture variability nor take into account factors of the perceived speech diversity (i.e., recording conditions, prosody, styles, and accents). Being able to measure speech diversity would also help detect mode collapse when finetuning models (Kirk et al., 2023), selecting training & inference hyperparameters, and optimizing for human feedback (Cideron et al., 2024). Moreover, such measurements are vital when building synthetic data or mixing existing natural datasets.
@@ 2 Background
Different diversity metrics have been proposed for evaluating generative models in computer vision (Heusel et al., 2017; Naeem et al., 2020; Sajjadi et al., 2018; Kynkäänniemi et al., 2019; Jalali et al., 2023) and natural language processing (Zhang et al., 2020; Li et al., 2016; Zhu et al., 2018; Shen et al., 2019; Dušek et al., 2019; Shu et al., 2019). However these metrics are either computed on raw inputs for text (Zhu et al., 2018; Shen et al., 2019; Alihosseini et al., 2019) or based on underlying representations specific to text and images (Heusel et al., 2017; Cífka et al., 2018; Tevet and Berant, 2021) which makes them unsuitable for evaluating acoustic diversity in generative speech models.
@@ 3 Learning Representations
We take a two-steps approach for measuring diversity of a set of speech utterances. First, we encode raw waveform audio in a compact representation and, afterwards, we estimate the diversity of a set of utterances in this embedding space. Hence, it is vital that the information about the relevant properties of speech is preserved in the embedding space, while the irrelevant details are omitted. We calculate embeddings in a hierarchical way: at first, we map audio to generic embeddings and then apply lightweight specialized projection models that are tailored to capture a specific facet. As a result, our approach allows us to get reliable per-facet diversity metrics. More details can be found in Section 6.1.
SpeechSim We aim to learn speech representations where acoustically similar audio segments are close to each other in the embedding space. To do so, we train a self-supervised contrastive model with positive pairs coming from non-overlapping fixed-length chunks of the same speech utterance. The main idea being that such speech segments are likely to be acoustically similar. We refer to this model as SpeechSim. SpeechSim provides a speech representation that are sensitive to changes in acoustic diversity, however, it has all types of acoustic properties mixed up together. Indeed, our experiments in Section 7 show that the vanilla SpeechSim embeddings are strongly biased towards speaker voice. We therefore learn projection models (heads) on top of SpeechSim embeddings to extract different types of acoustic information.
@@ 5.1 Our Approach
In the following, we describe the general template for building a series of datasets of controllable diversity, using voice diversity as a reference. We fix the sample size a priori and define voice diversity as the number of distinct speakers in a set of examples. To build one dataset, we sample examples from a natural dataset such that we obtain the required number of speakers and have every speaker represented by the same number of utterances. We create a series of datasets by varying the number of distinct speakers. We also ensure that other facets do not differ from one dataset to another as much as possible so that we make sure we indeed test speaker voice diversity.
@@ 5.2 Evaluation data
For each facet and each diversity level, we randomly sample 100 sets using different seeds.
@@
We build evaluation (validation) sets from LibriTTS (Zen et al., 2019) test-clean & test-other (dev-clean & dev-other). Each set contains 200 utterances of a single gender (either all male voices or all female voices) to test whether a particular metric is sensitive to the voice diversity independently from the gender diversity. We vary the number of distinct voices in $\{5,10,15,20,25,33\}$.
@@
We build evaluation (validation) sets from test (dev) sets of VCTK (Veaux et al., 2016). Accent classes are common between train, dev, and tests sets. Again, we built two types of samples (1) with equal number of different speaker voices and (2) with the minimum number of different speaker voices for high-entropy datasets but all speaker voices— equally represented—for low-entropy datasets. Speakers in test and dev sets are distinct from those found in the training set.
@@
We build evaluation and validation sets by sampling 20% and 10% of AudioSet, respectively (Gemmeke et al., 2017). We select examples that have exactly two tags, one of them being “Speech”. We treat the second tag as a class of the background noise. Next, we construct a sample by randomly selecting 100 examples with varying number of distinct noise classes (1, 5, 10, 25, 50, or 100).
@@
6 Experimental Setup
Details about the training data and the baseline representations we use can be found in Appendix A and Appendix B respectively.
6.1 Training details
SpeechSim
Projection models
We train projection models on top of SpeechSim embeddings with. a standard contrastive loss (Radford et al., 2021). The positive pairs are constructed by taking two examples with the same labels (e.g., same emotion). Except for the gender projector, all models have 2 fully connected layer (of size 256 and 128) with a GELU activation (Hendrycks and Gimpel, 2016) in-between. In the case of the gender projector, we found that a 4-layers network works better (dimensions of 256, 256, 128 and 128). Dropout rate of 0.1 is applied to SpeechSim input representations. We trained the projection models using Adam (Kingma and Ba, 2014) and use a learning rate of 1e-4, and a batch size of 128. We also apply a weight decay of 1e-3 for projection models related to speaker voices, emotion and background noise and 1e-4 for other facets. We select hyperparameters using validation loss.
@@ 7.1 Variability in a single facet
From Table 1 we firstly notice that, across all columns, SpeechSim shows higher correlation scores than off-the-shelf embedding models and the models trained from scratch. Further voice specialization of SpeechSim (SpeechSim/Voice) results in similar scores. Finally, on comparing average cosine vs. Vendi Score aggregation, we observe that often Vendi Score shows higher correlations.
@@
In Table 3 we report average Spearman correlations for the tested representation models, when tested on EmoV and Expresso separately. From these results, we see that, generally, SpeechSim performs better than other generic representations across all configurations. Equally, the specialized projection of SpeechSim for the emotion facet, SpeechSim/Emotion performs best across all representations. As with voice diversity, Vendi Score gets higher correlations than average cosine dissimilarity.
To make sure that our models can measure diversity of emotions beyond that is covered by the (limited) label sets of EmoV and Expresso, we run an experiment where the projection model is trained on one dataset and tested on another (the only common label is “neutral”). Table 3 shows that SpeechSim finetuned on Expresso does transfer to EmoV classes as it obtains a high Spearman correlation. However, SpeechSim finetuned on EmoV gets a Spearman correlation equal to 0.321. Digging deeper, we additionally compute the correlation of SpeechSim/Emotion-EmoV scores where we average all scores (obtained with different seeds) so that we obtain a single score for each diversity level. In this, SpeechSim/Emotion-Emov gets a correlation of 0.964 which means that SpeechSim finetuned on EmoV generalizes well on average.
@@
We report our results in Table 5. We see that generally all representations have somewhat low levels of correlation and high variability across seeds probably due to the fact that AudioSet is pretty noisy and multiple labels can be found in a single example. That said, SpeechSim/Noise combined with Vendi Score obtains the second highest correlation, following HuBERT.
We additionally compute the correlation of the SpeechSim/Noise scores where we average all scores (obtained with different seeds) so that we obtain a single score for each diversity level. In this setup, SpeechSim/Noise gets a Spearman correlation of 1.00, indicating that it works very well on average.
All in all, Tables 1 to 5 show that SpeechSim generic representations achieve higher correlation scores than off-the-shelf embedding models while having lower latency, this motivates us to rely on it as a basis for per-facet models.
@@ 7.2 Are facets mixed within metrics?
We run a similar experiment for accents and emotions and report it in Appendix D. We observe similar findings, raw SpeechSim is highly influenced by Speaker voice variations while specialized facets are unaffected by changes in voices.
Overall, the experiments reported in this Section and Appendix D support the necessity of introducing per-facet projection models. While SpeechSim embeddings might provide reasonably good correlations with the ground-truth diversity when only one facet is changed, they implicitly mix all facets into a single score. We find this inconvenient, as (a) changes in one facet can mask changes in another, (b) the user might want to control the trade-off explicitly. Hence, per-facet models are preferrable.
8 Comparing Generative Models
We start by assessing whether the introduction of SoundStorm (Borsos et al., 2023b) to the AudioLM framework Borsos et al. (2023a) lead to any hidden penalties in acoustic diversity (Section 8.1) before comparing acoustic diversity of off-the-shelf TTS models (Section 8.2). Additional studies on the effect of the Best-K sampling technique used by Kharitonov et al. (2023) on acoustic diversity and how temperature changes acoustic diversity of the synthesized speech can be found in Appendix C.1 and Appendix C.2. In this Section, we use the same terminology as Borsos et al. (2023a) where semantic tokens refer to tokens learned with a self-supervised masked language modeling objective (Chung et al., 2021a; Hsu et al., 2021) while acoustic tokens refer to tokens learned with a full reconstruction objective from a residual vector quantizer (Zeghidour et al., 2021).
8.1 Semantic-to-Acoustic Token Conversion: SoundStorm vs. AudioLM
In Table 7 we report, for each facet of the acoustic diversity, the ratio of utterances where AudioLM generated higher-diversity audio than SoundStorm. We see that SoundStorm produces audio that is more diverse in voices, emotions, accents, and background noise, while being behind in the gender diversity.
Voice Gender Emotion Accent Background noise
0.0 62.5 6.3 1.6 4.7
Table 7: Semantic-to-acoustic tokens resynthesis task. The proportion of sets of utterances (%) where the AudioLM stage 2 model generates samples that are more diverse than those generated by SoundStorm.
8.2 Acoustic diversity in Text-to-Speech systems
Next, we measure acoustic diversity of several publicly available TTS systems. Here, our goal is to show that MAD Speech brings a new dimension to the evaluation of real-world TTS systems. We compare Bark TTS [^2] https://github.com/suno-ai/bark, StyleTTS 2 (Li et al., 2024), Tortoise TTS (Betker, 2023a) and FastSpeech 2 (Ren et al., 2020). Details about these models can be found in Appendix E. As some of these systems are deterministic, we adopt a different setup and synthesize all transcripts in shuffled LibriTTS test-clean and calculate the average acoustic diversity across batches of audio of size 128. We report the results in Table 8.
Voice Gender Emotion Accent Back. Noise
Bark TTS 39.41 0.90 8.29 8.30 3.84
Tortoise TTS 30.94 0.92 8.73 8.17 3.17
StyleTTS 2 31.54 0.37 7.73 6.81 2.60
FastSpeech 2 19.33 0.29 6.74 5.63 2.42
We see that Bark and Tortoise TTS are the most diverse across all facets, likely due to training on more diverse internet data. Interestingly, Tortoise TTS tends to be more diverse than Bark in emotions while less or equally diverse in other facets. StyleTTS 2 is considerably less diverse in the gender facet — just as one would expect, given that the generation is restricted to female voices. FastSpeech 2 is the least acoustically diverse mainly due to being trained on a very homogeneous dataset.
All in all, the comparisons we carried out in this Section and Appendix C show that previously proposed changes lead to non-trivial and sometimes even non-monotonic changes in per-facet acoustic diversity. Equally, the sampling temperature also affects the resulting acoustic diversity. We also found that MAD Speech can be useful in detecting acoustic diversity trends in off-the-shelf TTS systems. Overall, we believe those findings highlight the necessity for including acoustic diversity metrics in the standard evaluations.
9 Conclusions & Future work
We introduced MAD Speech, a set of metrics for evaluating acoustic diversity in speech. MAD Speech metrics are calculated in two steps: (1) a set of utterances is mapped to an embedding space, (2) acoustic diversity is estimated by comparing the resulting embeddings with an aggregation function (cosine dissimilarity or Vendi Score). Calculating embeddings is done by firstly applying a generic representation model with a subsequent metric-specific projection. We focus on five facets of acoustic diversity: voices, gender, emotion, accents, and background noise.
Finally, we highlighted a necessity of controlling acoustic diversity when developing new models and approaches. To this end, we took some recent modelling improvements and demonstrated that they, in fact, affect acoustic diversity of speech.
Limitations
By building MAD Speech, we make it possible to holistically assess progress in generative speech and reveal possible hidden biases in the acoustic scenes. Yet, our metrics are machine-learned models themselves and are limited by the nature of the data we used to build them. Specifically, the public datasets we use are limited to English language. The labels in the data can be noisy and only represent a limited reflection of the world. We also made the choice to treat gender as a binary variable to be consistent with existing datasets, this can lead to minor measurement inaccuracies when facing ambiguously-gendered voices (Stoidis and Cavallaro, 2022).
Appendix A Training Data
Appendix B Baseline Representations
B.1 Generic representations
Wav2Vec-BERT
HuBERT
SoundStream
B.2 Specialized representations
Training from scratch
Appendix C Effect of decoding method and temperature on acoustic diversity
C.1 SPEAR-TTS: Effect of Best-of-K decoding
K Voice Gender Emotion Accent Background Noise
1 64 41 62 63 61
4 0 27 0 0 0
C.2 Text Synthesis: Influence of Temperature
T Voice Gender Emotion Accent Background Noise
0.7 12 33 21 18 29
0.9 45 21 40 40 36
Table 10: Count (out of 64) of instances where sets of utterances generated with $T\in\{0.7,0.9\}$ have higher diversity scores than with $T=0.8$.
Appendix D Additional experiments on testing whether facets are mixed up together
D.1 Emotions and Voices
@@
We examine the interactions between emotion and voice diversity within metrics as explained in Section 7.2. The number of speakers is 1 or 4 depending on the diversity level of the set of utterances. For sets with high entropy (i.e. high diversity), there is an unique speaker voice while there are 4 speakers for sets with low entropy (i.e. low diversity). We report the results in Table 11. We observe that SpeechSim/Emotion is unaffected by the changes in the number of speakers, while SpeechSim/Voice follows the speaker diversity and hence it has a negative correlation with the emotion diversity score. SpeechSim focuses on the changes in the speaker diversity, thus having a negative correlation with the gender diversity.
D.2 Varying accents and voices in opposite directions
@@
In this experiment, we use 7 voices for sets with high entropy and 30 voices for sets with low entropy. The results are reported in Table 12. We see that the projected metric is unaffected by changes in voices, while the vanilla version of SpeechSim is focused on the speaker facet.
@@
Appendix E Acoustic diversity in Text-to-Speech systems - Model details